OpenStudy (anonymous):

FLVS HELP PLEASE Help please !!! Part 1. Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses. f(x)= x+a b g(x)=cx−d Part 2. Show your work to prove that the inverse of f(x) is g(x). Part 3. Show your work to evaluate g(f(x)). Part 4. Graph your two functions on a coordinate plane. Include a table of values for each function. Include 5 values for each function. Graph the line y = x on the same graph.

OpenStudy (anonymous):

What unit is this?

OpenStudy (anonymous):

its the 5.08 project -.-

OpenStudy (anonymous):

have you done it before ?

OpenStudy (anonymous):

I did this through connections, I remember doing it. I'll see if I still have the file so I can help you.

OpenStudy (anonymous):

thank you !!! go in the gradebook tab

OpenStudy (anonymous):

@Maria977 you there?

OpenStudy (anonymous):

@helder_edwin

OpenStudy (anonymous):

@PurplePunch

OpenStudy (anonymous):

I couldn't find the file, as I took a multiple choice test on this. But I did manage to find the lesson.

OpenStudy (anonymous):

But, do you have any idea on what you want a, b and c to be?

OpenStudy (anonymous):

@Maria977 no i dont know anything about functions, my grade went down to a C just because of this section

OpenStudy (anonymous):

(functions)

OpenStudy (anonymous):

Okay, how about we do f(x) = 3x -5 ?

OpenStudy (anonymous):

ok ... ? im going by what you say cuz i dont understand it anyway

OpenStudy (anonymous):

but ok

OpenStudy (anonymous):

@Maria977 are you typing still ?

OpenStudy (anonymous):

well, to get g(x) you just find the inverse of f(x) like this: f(x) = 3x -5 Now change f(x) for y= 3x -5 Switch the x and y x=3y -5 Now isolate the y by adding five to both sides x + 5 =3y Now divide both sides by 3 \[ \frac{ x +5 = 3y }{ 3 }\]\[y = \frac{ x +5 }{ 3 }\]

OpenStudy (anonymous):

this is just part 1 right ?

OpenStudy (anonymous):

For part one, you just write the functions, and then for part 2, you show the work that I showed to find the inverse.

OpenStudy (anonymous):

So you basically gave me the full answer for all of task 1 ?

OpenStudy (anonymous):

@Maria977

OpenStudy (anonymous):

@Maria977 is that it...? can you help me with task 2 ?

OpenStudy (anonymous):

do you know what composition of functions is?

OpenStudy (anonymous):

It looks familiar but no, ive read the lesson twice and watched youtube videos and i just cannot understand it

OpenStudy (anonymous):

so no..

OpenStudy (anonymous):

Ok, so you question is asking for g(f(x)). I am not that sure on this, but I think it should be your f(x) function replaces the x part in your g(x) function.

OpenStudy (anonymous):

thats part 3 right

OpenStudy (anonymous):

What I mean is \[g(f(x))= \frac{ (3x-5) +5 }{ 3 }\] I think it looks so difficult because of the f(x) function I chose

OpenStudy (anonymous):

no its not what you chose, i just literally have never understood algebra, its my second time taking it. the equation you just put, is that the answer for part 3 ?

OpenStudy (anonymous):

you see what I mean by replacing the x in g(x), right? I think we should try this with a different f(x) function, yes?

OpenStudy (anonymous):

no i dont see because i dont understand and sure whatever you say, if you could just please tell me what is what part

OpenStudy (anonymous):

ill try to understand but i most probably will not with more and more numbers and information just please help me with answers or equations

OpenStudy (anonymous):

ill still try

OpenStudy (anonymous):

the last part I showed you, g(f(x)) is part 3

OpenStudy (anonymous):

so is this all i need to submit for task 1 ? Task 1 well, to get g(x) you just find the inverse of f(x) like this: f(x) = 3x -5 Now change f(x) for y= 3x -5 Switch the x and y x=3y -5 Now isolate the y by adding five to both sides x + 5 =3y Now divide both sides by 3 x+5=3y / 3 y=x+5/ 3 g(f(x))=(3x-5)+5 /3

OpenStudy (anonymous):

No, for part one submit F(x) = 3x-5 (g(x) = \[\frac{ x+5 }{ 3 }\]

OpenStudy (anonymous):

what about the y=x+5/3

OpenStudy (anonymous):

what about all the steps you showed me before..

OpenStudy (anonymous):

after you finish finding the inverse, you are supposed to replace the y with g(x). It is still the same function, but make sure you put g(x) instead of y

OpenStudy (anonymous):

The steps are for part 2

OpenStudy (anonymous):

wait so

OpenStudy (anonymous):

F(x) = 3x-5 (g(x) = x+5/3 is the answer for JUST part 1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and then part 2 is f(x) = 3x -5 Now change f(x) for y= 3x -5 Switch the x and y x=3y -5 Now isolate the y by adding five to both sides x + 5 =3y Now divide both sides by 3 x+5=3y / 3 y=x+5/ 3

OpenStudy (anonymous):

yes, because part two ask you to prove how g(x) is the inverse of f(x), s you show the steps as proof

OpenStudy (anonymous):

and part 3 ... ??

OpenStudy (anonymous):

part 3 is the g(f(x)) function I showed you

OpenStudy (anonymous):

which is ?

OpenStudy (anonymous):

i got confused

OpenStudy (anonymous):

look back towards the beginning of our conversation, and find the g(f(x)) function

OpenStudy (anonymous):

did you find it ? If not it is: \[g(f(x))=\frac{ (3x-5)+5 }{ 3}\]

OpenStudy (anonymous):

@Maria977 so that is part 3 ? isnt that part 1 ?

OpenStudy (anonymous):

@Maria977 oh never mind, so thats it, i have all 3 parts ?

OpenStudy (anonymous):

Yep, do you know how to get part four?

OpenStudy (anonymous):

no because its a graph but i dont know how to graph it on a computer or how to type in a graph

OpenStudy (anonymous):

okay, wait a sec.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The blue is f(x) and the red is g(x) ok

OpenStudy (anonymous):

so i attach this to the assingment ?

OpenStudy (anonymous):

for part 4 ?

OpenStudy (anonymous):

Yes, attach it as part of the assignment. Also, for part 3, make sure to include some evidence. For example, write that f(x) replaces the x in the function g(x) and also include the g(f(x)) function.

OpenStudy (anonymous):

thank you so much!!! you saved me

OpenStudy (anonymous):

@Maria977